OpenStudy (anonymous):

How do you solve the equation by using the square root property for the following problem; 3x^2=5

OpenStudy (psymon):

The best you could do is get x^2 by itself then square root both sides. Just know that when you take the square root of both sides, you must put plus or minus. If that makes sense.

OpenStudy (anonymous):

kinda??? I was using a Calculator, and it gives me a large number with a decimal, with a minus and a plus in front of it, although it's not giving me the what I'm looking for...

OpenStudy (psymon):

Ah. Well, let's get x^2 by itself first by dividing both sides by 3: \[x ^{2}=\frac{ 5 }{ 3 } \] Then we would take the square root of both sides:

OpenStudy (psymon):

\[\sqrt{x ^{2}}=\sqrt{\frac{ 5 }{ 3 }}->x=\pm \sqrt{\frac{ 5 }{ 3 }}\]

OpenStudy (anonymous):

\[3x ^{2}=5 \] \[x ^{2}= \frac{ 5 }{ 3 }\] \[x = +- \sqrt{\frac{ 5 }{ 3 }}\]

OpenStudy (anonymous):

got it thx -15/3,15/3

OpenStudy (psymon):

Just make sure to put those 15's as square roots :P

OpenStudy (anonymous):

gotcha, thx... I appreciate your help.

OpenStudy (psymon):

No problem :3